OpenStudy (anonymous):

If f(x)=x+5 and h(x)=4x-5, find a function g such that g•f=h

OpenStudy (anonymous):

like solving \(3\times g=10\) you would get \(g=\frac{10}{3}\) in this case you are solving \[(x+5)\times g(x)=4x-5\]solve by division

OpenStudy (anonymous):

Wait that's not an option do you think it would be 5x-4??

OpenStudy (anonymous):

@satellite73 I'm a bit confused that's not an option